# Splittable Spanning Trees and Balanced Forests in Dense Random Graphs

### v2

@@ 1. Introduction
When used for redistricting, random weighted equitable partitions are sampled to form an ensemble against which proposed maps can be tested for political gerrymandering. Each vertex of the graph represents a small indivisible region, or precinct, and the components of a partition represent the districts from which representatives are elected. This methodology has been used for court cases and various amicus briefs [charikar2022complexity, zhao2022, DukeReCom1]. While the requirements for acceptable political maps vary state to state, mathematicians have generally agreed that the spanning tree weight is a desirable baseline because spanning trees are easy to generate and they typically produce maps with compact districts [procacciaTuckerFoltz2021compactness]; moreover, when other weightings are preferable, parallel tempering has been successfully used to interpolate between the spanning tree weight and other useful metrics [DukeReCom1, DukeReCom2]. Uniformly generating balanced forests is also desirable in the contexts of clustering in biological and social networks [camilus2012, junker] and image decomposition [fortunato2010].
1.1. Algorithms for sampling balanced forests
@@
A tree is $k$-splittable if there exist k1k-1 edges that, when removed, create a graph whose $k$ connected components all have the same number of vertices.
@@
The most widely used algorithm for sampling balanced forests is ReCom (for recombination) [recom] and its variants, which provably converge to the distribution defined by spanning tree weight [revrecom, DukeReCom1] [^1] For simplicity of exposition, we refer to this group of algorithms simply as “ReCom,” although our results rely on any of the modified versions which have provably correct stationary distributions [revrecom, DukeReCom1]. ReCom uses Markov chains that walk on balanced forests (or weighted partitions) by iteratively choosing two adjacent pieces, finding a spanning tree on the vertices in the union of these piece, and then partitioning this joint region into two equal sized pieces by splitting the spanning tree in two, if possible. When a tree cannot be equitably split by removing a single edge, the move is rejected and the Markov chain proceeds to propose a new recombination move.
@@
For this approach to be efficient and reliable, it is necessary that the recombination move connects the state space of partitions, the chain is rapidly mixing, and subgraphs of $G$ contain splittable spanning trees with at least an inverse polynomial probability. Unfortunately, it has been shown that there are potentially issues with each of these three requirements. In particular, no graphs except the complete graph are known to be connected under perfectly balanced ReCom moves. Moreover, Charikar et al. gave a family of graphs where strictly following ReCom moves will be slowly mixing [charikar2022complexity]. Natural modifications to ReCom have been proposed to increase its connectivity and potentially its mixing rate by allowing some slack on the piece sizes. Yet, even with this allowance, the state space has only been shown to be connected in special cases [charikar2022complexity, cannontri, akitaya2022reconfiguration_recombination].
@@
Another family of sampling algorithms work by generating uniform spanning trees on all of $G$ (using, for example, the loop erased random walk (LERW) algorithm of Wilson [wilson]) and checking whether each tree is $k$-splittable, which can be done efficiently if $k$ is constant. Note that any splittable tree has a unique set of k1k-1 edges whose removal leaves a balanced forest (see Lemma 2.1), so rejection sampling will generate such forests uniformly. If the tree is not splittable, it is resampled. This algorithm produces samples from exactly the right conditional distribution, but the running time depends on the rejection rate.
@@
Markov chains that walk on the set of (not necessarily balanced) forests with $k$ components also exist, such as the up-down walk [charikar2022complexity]. This chain has a $O(n^{2}\log(n))$ mixing time on all graphs and is known to be ergodic, but again we need to employ rejection sampling whenever the forests are imbalanced. The remaining balanced forests will be generated almost uniformly and all other forests will be rejected and resampled, so the overall algorithm will be fast only if the rejection rate is not too large. Note that the probability that a random tree is splittable is polynomially related to the likelihood a random forest is balanced (see Lemma 2.2), so either both the LERW and up/down walk approaches to sampling balanced forests are efficient or neither will be.
@@
As all of these algorithms rely on the probability that random spanning trees are splittable, it is worthwhile to characterize which classes of graphs have this property. Charikar et al. [charikar2022complexity] conjectured that many graphs are splittable with inverse polynomial probability, including rectangular regions of $\mathbb{Z}^{2}$. Recently, Cannon, Pegden and Tucker-Foltz [cannon2024grid] gave a stunning proof of this conjecture for planar grid graphs and grid-like graphs. Planarity is critical to their argument because they rely on the dual of the loop erased random walk to bound the likelihood of specific families of forests being generated to show that these already occur with at least $1/n^{2}$ probability, suppressing dependencies on $k$.
@@
Little is known about sufficient conditions to ensure splittability in other classes of graphs. Akitaya et. al. [akitaya2022reconfiguration_recombination] have conjectured that Hamiltonicity is sufficient, and Cannon et. al. [cannontri] have conjectured that it is sufficient for a graph to be an induced subgraph of the triangular lattice. Little is known as well about sufficient conditions to ensure fast mixing of ReCom. It has been conjectured that it is necessary to allow slack on the piece sizes to guarantee connectivity of the state space [akitaya2022reconfiguration_recombination, cannontri], or to avoid slow mixing [charikar2022complexity].
1.2. Results
@@
While it is tempting to believe the conventional wisdom that there are simple conditions on graphs that are sufficient to ensure their splittability with high probability, such as Hamiltonicity, sparsity, and planarity (e.g., [akitaya2022reconfiguration_recombination, cannontri, charikar2022complexity]), we give counterexamples to each of these conjectures. In particular, in Theorem 3, Section 3, we present an infinite family of graphs on $3n$ vertices that are induced Hamiltonian subgraphs on the triangular lattice, where the only equitable partitions of the graph into three connected components are the $n$ cyclical partitions of the Hamiltonian cycle at i,i+n,i+2n,i,\ i+n,\ i+2n, for $0\leq i\leq n-1$. However, the probability of a random spanning tree being splittable on this family is exponentially small. This implies that all of the algorithms for generating balanced partitions mentioned above, including ReCom, LERWs, and the up-down walk, all provably require exponential time.
@@
We have the following theorem.
Theorem 1.
Finally, we prove that a polynomial fraction of the set of k-component forests on dense regular graphs will be balanced, for any constant k. This complements the work of Cannon et al. [cannon2024grid], who established a similar result for the m x n grid and a related class of “grid-like” graphs.
Theorem 2.
Note that Cannon et al. [cannon2024grid] showed the proportion of spanning trees that are splittable on the $\sqrt{n}\times\sqrt{n}$ grid scales by at least $n^{2-2k}$, so these bounds on splittability for dense graphs are much more favorable.
These are the first results showing any dense graphs are splittable with large enough probability to use rejection sampling, as needed by the simple algorithm, the up-down walk, and ReCom. Interestingly, the results on $G(n,m)$ are more straightforward to prove. We then derive the matching results for $G(n,p)$, rather than the other way around.
What makes analysis of the splittability of dense graphs especially challenging, and even somewhat surprising, is that spanning trees that appear with higher probability than others are less likely to be splittable. While any two spanning trees are equally likely to appear in a random graph, one tree may be more likely to appear in graphs that contain fewer spanning trees, thus making it more likely to be selected at random. For example, a star with three leaves is more likely than a path of three edges to be selected as a random spanning tree of a random graph on four vertices. Alon et al. [alon20] have shown that the diameter of random spanning trees of dense random graphs grow as $\sqrt{n}$, as was previously shown by Szekeres [szekeres83] for the complete graph. Trees with smaller diameter would seem less likely to be splittable than those with large diameter, so inverse polynomial bounds on the splittability of uniform spanning trees on dense graphs is more suprising than on sparse graphs such as the grid, where the diameter of a uniform spanning tree is linear.
Another difficulty is that balanced forests tend not to be the mode of a histogram showing the probabilities of splits of various sizes, even when $k=2$. In fact, when on the complete graph, balanced splits contain the fewest forests out of all possible partition sizes. (See Figure 2 in Section 4.) This means that standard techniques showing that this distribution is log-concave or unimodal will certainly fail and the calculations need to be more precise.
2. Preliminaries
@@
2.1. The spanning tree weight and splittability
@@
Definition 2.1.
The Spanning Tree Weight of a connected partition $P$ of a graph $G$ is the product of the number of spanning trees in each piece.
@@
If $F_{2}$ contains no edges in $S$, all of $S$ is connected in $T-F_{2}$. Then to avoid the connected component containing $S$ having more than $n$ vertices, $F_{2}$ must contain $e$, as otherwise, $S$ is connected and connected to $e$ in $T-F_{2}$. Then both $F_{1}$ and $F_{2}$ contain $e$, and removing $e$ splits $T$ into a connected component with $n$ vertices and one with (k1)n(k-1)n vertices. Then both $F_{1}$ and $F_{2}$ contain subsets of k2k-2 edges splitting this other subgraph into pieces of size $n$, so by induction, they contain the same edges. Then $F$ is unique. ∎
Lemma 2.2.
@@
    There exists a polynomial $p_{k}(n)$ such that for any $G=(V,E)\in\mathcal{G}$, a random spanning tree of $G$ is $k$-splittable with probability at least $\frac{1}{p_{k}(|V|)}$
  2. (2)
@@
    There exists a polynomial $q_{k}(n)$ such that for any $G=(V,E)\in\mathcal{G}$, a random $k$-partition of $G$ weighted by the spanning tree weight is balanced with probability at least $\frac{1}{q_{k}(|V|)}$.
Proof.
@@
We note that the probability that a random spanning tree is balanced is the ratio of the number of $k$-splittable spanning trees to the number of spanning trees of a graph, and that the probability that a random k-partition weighted by the spanning tree weight is balanced is the ratio of the total weight of all balanced $k$-partitions to the total weight of all $k$-partitions. Then this is a direct consequence of the next two lemmas. A similar result is proven in [cannon2024grid]. ∎
Lemma 2.3.
@@
For any graph $G$ with $n$ vertices, the ratio between the total weight of all $k$-partitions of $G$ and the number of spanning trees of $G$ is between $n^{-2(k-1)}$ and $n^{2(k-1)}$.
Proof.
@@
The weight of a partition $P$ under the spanning tree weight is the product of the number of spanning trees in each piece. Then this weight is equal to the number of spanning forests whose connected components are the pieces of $P$. Then consider the space of spanning forests in $G$ with $k$ connected components.
@@
We see that each such forest can, by adding k1k-1 edges of the components, be made into a spanning tree, and since every spanning tree can be split into $k$ pieces by arbitrarily removing k1k-1 edges, every spanning tree can be made this way. Since there are at most $n^{2}$ edges in $G$, the number of spanning trees is at most $n^{2(k-1)}$ times the number of such forests. Similarly, every such forest can be made by removing k1k-1 edges from a spanning tree, the number of such forests is at most $n^{2(k-1)}$ times the number of spanning trees (this bound can be reduced to $n^{k-1}$ by the structure of the tree). ∎
Lemma 2.4.
@@
For any graph $G$ with $n$ vertices, the ratio of the total weight of all balanced $k$-partitions of $G$ and the number $k$-splittable of spanning trees of $G$ is between $n^{-2(k-1)}$ and $n^{2(k-1)}$.
Proof.
@@
By Lemma 2.1, each splittable tree uniquely defines a $k$-balanced spanning forest, so the number of them is at most the number of spanning trees. Further, each balanced tree can be defined by a balanced forest and k1k-1 edges, so there are at most $n^{2(k-1)}$ times as many spanning trees as balanced forests. ∎
2.2. Algorithms for generating balanced forests
It will be useful to formalize the three main algorithms for sampling balanced partitions.
@@ 2.2. Algorithms for generating balanced forests
First, the simplest algorithm is splitting a full spanning tree (see Algorithm 1). This algorithm will run a single stage and either return a balanced partition sampled uniformly according to the spanning tree weight, or fail. When using it in practice, the algorithm will be run until it succeeds in finding a balanced $k$-partition. It is described in [cannon2024grid]. The algorithm first generates a random spanning tree. Then it generates a partition by splitting the tree into $k$ equal pieces, if that is possible. Then it accepts the partition with probability equal to the inverse of the number of spanning trees that would generate that partition. The algorithm finds k1k-1 edges in the spanning tree to split it equitably as follows. If k1k-1 such edges exist, they are unique, according to Lemma 2.1. Each edge, when removed by itself from the tree, splits the tree into two pieces whose sizes are multiples of $n$, and this property uniquely characterizes these k1k-1 edges.
@@
6:  remove the edges in $K$ from the tree to create a forest
@@
Next, we will examine two Markov-chain based algorithms. The up-down forest walk, described in [charikar2022complexity], moves among forests with NkN-k edges (see Algorithm 2). The number of times $t$ that the loop is iterated depends on the mixing time, which was shown to be $O(n^{2}\log(n))$ in the worst case [charikar2022complexity].
@@
The final algorithm we describe is Reversible ReCom as described in [revrecom] (see Algorithm 3). This is also a Markov Chain algorithm that only moves between balanced partitions, so always returns a valid balanced partition, however its mixing time is unknown and sometimes exponential.
@@
All of the listed Markov Chain algorithms are known to converge to a stationary distribution that selects each forest with equal probability, including both reversible ReCom and Forest ReCom[charikar2022complexity][DukeReCom1][revrecom]. However, ReCom Markov chains may not be irreducible [akitaya2022reconfiguration_recombination], so may have multiple stationary distributions depending on the starting partition. Moreover, guarantees on the mixing times of ReCom are unknown, with some cases known to require exponential time[charikar2022complexity].
3. Challenges when allowing slack

### v3

@@ 1 Introduction
When used for redistricting, random weighted equitable partitions are sampled to form an ensemble against which proposed maps can be tested for political gerrymandering. Each vertex of the graph represents a small indivisible region, or precinct, and the components of a partition represent the districts from which representatives are elected. This methodology has been used for court cases and various amicus briefs [6, 12, 21]. While the requirements for acceptable political maps vary state to state, mathematicians have generally agreed that the spanning tree weight is a desirable baseline because spanning trees are easy to generate and they typically produce maps with compact districts [17]; moreover, when other weightings are preferable, parallel tempering has been successfully used to interpolate between the spanning tree weight and other useful metrics [5, 6]. Uniformly generating balanced forests is also desirable in the contexts of clustering in biological and social networks [7, 16] and image decomposition [14].
1.1 Algorithms for sampling balanced forests
@@
A tree is $k$-splittable if it contains k1k-1 edges that, when removed, create a graph whose $k$ connected components all have the same number of vertices.
@@
The most widely used algorithm for sampling balanced forests is ReCom (for recombination) [13] and its variants, which provably converge to the distribution defined by spanning tree weight [6, 8] [^1] For simplicity of exposition, we refer to this group of algorithms simply as “ReCom,” although our results rely on any of the modified versions which have provably correct stationary distributions [6, 8]. ReCom uses Markov chains that walk on balanced forests (or weighted partitions) by iteratively choosing two adjacent pieces, finding a spanning tree on the vertices in the union of these pieces, and then partitioning this joint region into two equal sized pieces by splitting the spanning tree in two, if possible. When a tree cannot be equitably split by removing a single edge, the move is rejected and the Markov chain proceeds to propose a new recombination move.
@@
For this approach to be efficient and reliable, it is necessary that the recombination move connects the state space of partitions, the chain is rapidly mixing, and subgraphs of $G$ contain splittable spanning trees with at least an inverse polynomial probability. Unfortunately, it has been shown that there are potentially issues with each of these three requirements. In particular, no graphs except the complete graph are known to be connected under perfectly balanced ReCom moves. Moreover, Charikar et al. gave a family of graphs where strictly following ReCom moves will be slowly mixing [12]. Natural modifications to ReCom have been proposed to increase its connectivity and potentially its mixing rate by allowing some slack on the piece sizes. Yet, even with this allowance, the state space has only been shown to be connected in special cases [3, 10, 12].
@@
Another family of sampling algorithms work by generating uniform spanning trees on all of $G$ (using, for example, the loop erased random walk (LERW) algorithm of Wilson [20]) and checking whether each tree is $k$-splittable, which can be done in $O(n)$ time by a depth-first search. Note that any splittable tree has a unique set of k1k-1 edges whose removal leaves a balanced forest (see Lemma 1), so rejection sampling will generate such forests uniformly. If the tree is not splittable, it is resampled. This algorithm produces samples from exactly the right conditional distribution, but the running time depends on the rejection rate.
@@
Markov chains that walk on the set of (not necessarily balanced) forests with $k$ components also exist, such as the up-down walk [12]. This chain has a $O(n^{2}\log(n))$ mixing time on all graphs and is known to be ergodic, but again we need to employ rejection sampling whenever the forests are imbalanced. The remaining balanced forests will be generated almost uniformly and all other forests will be rejected and resampled, so the overall algorithm will be fast only if the rejection rate is not too large. Note that the probability that a random tree is splittable is polynomially related to the likelihood a random forest is balanced (see Lemma 2), so either both the LERW and up-down walk approaches to sampling balanced forests are efficient or neither will be.
@@
As all of these algorithms rely on the probability that random spanning trees are splittable, it is worthwhile to characterize which classes of graphs have this property. Charikar et al. [12] conjectured that many graphs are splittable with inverse polynomial probability, including rectangular regions of $\mathbb{Z}^{2}$. Recently, Cannon, Pegden and Tucker-Foltz [9] gave a stunning proof of this conjecture for planar grid graphs and grid-like graphs. Planarity is critical to their argument because they rely on the dual of the loop erased random walk to bound the likelihood of specific families of forests being generated to show that these already occur with at least $1/n^{2}$ probability, suppressing dependencies on $k$.
@@
Little is known about sufficient conditions to ensure splittability in other classes of graphs. Akitaya et al. [3] have conjectured that Hamiltonicity is sufficient, and Cannon et al. [10] have conjectured that it is sufficient for a graph to be an induced subgraph of the triangular lattice. Little is known as well about sufficient conditions to ensure fast mixing of ReCom. It has been conjectured that it is necessary to allow slack on the piece sizes to guarantee connectivity of the state space [3, 10], or to avoid slow mixing [12].
1.2 Obstacles to splittability
@@
While it is tempting to believe the conventional wisdom that there are simple conditions on graphs that are sufficient to ensure their splittability with high probability, such as Hamiltonicity, sparsity, and planarity (e.g., [3, 10, 12]), we give counterexamples to each of these conjectures. In particular, in Theorem 3.1 (Section 3), we present an infinite family of graphs on $3n$ vertices that are induced Hamiltonian subgraphs on the triangular lattice, where the only equitable partitions of the graph into three connected components are the $n$ cyclical partitions of the Hamiltonian cycle at i,i+n,i+2n,i,\ i+n,\ i+2n, for $0\leq i\leq n-1$. However, the probability of a random spanning tree being splittable on this family is exponentially small. This implies that all of the algorithms for generating balanced partitions mentioned above, including ReCom, LERWs, and the up-down walk, all provably require exponential time.
@@
1.3 Splittability of dense random graphs
We prove that dense graphs, on the other hand, do enjoy high splittability probability. We show that if a graph is sufficiently dense, it will be splittable with high enough probability that we can efficiently generate balanced partitions using either the LERW or the up-down chain. First, we observe that the mixing time of ReCom on $K_{kn}$ is proportional to the inverse of this probability of being splittable. To see this, we can consider a “bin-mixing” Markov chain on $k$ bins of $n$ balls each, where a move consists of combining $2$ bins and randomly mixing the balls between them such that each bin ends up with $n$ balls. We see that when running ReCom on the complete graph, each move where it finds a $2$-splittable spanning tree is equivalent to a move on the bin-mixing Markov chain, and each one where it does not is a move where it does nothing. ReCom on this graph can be considered as a lazy version of the bin-mixing Markov chain. If the bin-mixing Markov chain mixes in time $M$, then ReCom mixes in time proportional to $M$ divided by the probability that a random spanning tree of $K_{2n}$ is $2$-splittable.
More precisely, we show in Theorem 4.1 that for any constant $k$, the probability that the complete graph $K_{kn}$ is $k$-splittable is
@@ 1.3 Splittability of dense random graphs
Furthermore, in Theorem 4.2 we prove our main theorem, showing that a polynomial fraction of the set of k-component forests on dense regular graphs will be balanced, for any constant k. Specifically, we show that if $N=nk$ and $G$ is a random connected graph chosen from $G(N,p)$ with $p=\Omega(1/\log(n))$, then the probability that a random spanning tree is k-splittable is at least
[…]
The same holds for $G(n,m)$ for random graphs of comparable expected density. This complements the work of Cannon et al. [9], who established a similar result for the m x n grid and a related class of “grid-like” graphs. Note that Cannon et al. [9] showed the proportion of spanning trees that are splittable on the $\sqrt{n}\times\sqrt{n}$ grid scales by at least $n^{2-2k}$, so these new bounds on splittability for dense graphs are much more favorable.
These are the first results showing any dense graphs are splittable with sufficiently large probability to use rejection sampling, which is a requirement for each of the simple algorithm, the up-down walk, and ReCom. Interestingly, the results on $G(n,m)$ are more straightforward to prove. We then derive the corresponding results for $G(n,p)$ from those, rather than the other way around.
One of the challenges to bounding splittability on dense graphs is that the spanning trees that appear with higher probability than others are actually less likely to be splittable. While all spanning trees are equally likely to appear in a random graph, some trees may be more likely to appear in graphs that contain fewer spanning trees, making it more likely to be selected. For example, a star with three leaves is more likely than a path of three edges to be selected as a random spanning tree of a random graph on four vertices. In fact, Alon et al. [4] have shown that the diameter of random spanning trees of dense random graphs grow as $\sqrt{n}$, as was previously shown by Szekeres [19] for the complete graph. Trees with smaller diameter would seem less likely to be splittable than those with large diameter, so inverse polynomial bounds on the splittability of uniform spanning trees on dense graphs may be more surprising than similar results on sparse graphs such as the grid, where the diameter of a uniform spanning tree is linear.
2 Preliminaries
@@
2.1 The spanning tree weight and splittability
@@
Definition 2
The Spanning Tree Weight of a partition $P$ of a graph $G$ is the product of the number of spanning trees in each piece.
@@
If $F_{2}$ contains no edges in $S$, all of $S$ is connected in $T-F_{2}$. Then to avoid the connected component containing $S$ having more than $n$ vertices, $F_{2}$ must contain $e$, as otherwise, $S$ is connected and connected to $e$ in $T-F_{2}$. Then both $F_{1}$ and $F_{2}$ contain $e$, and removing $e$ splits $T$ into a connected component with $n$ vertices and one with (k1)n(k-1)n vertices. Then both $F_{1}$ and $F_{2}$ contain subsets of k2k-2 edges splitting this other subgraph into pieces of size $n$, so by induction, they contain the same edges. Then $F$ is unique.
It will be convenient to talk about both balanced forests and splittable trees. The following establishes their close relationship. A similar result appears in [9].
Lemma 2
@@
    There exists a polynomial $p_{k}(\cdot)$ such that, for any $G=(V,E)\in\mathcal{G}$, a random spanning tree of $G$ is $k$-splittable with probability at least ${1}/{p_{k}(|V|)}.$
  2. 2.
@@
    There exists a polynomial $q_{k}(\cdot)$ such that, for any $G=(V,E)\in\mathcal{G}$, a random $k$-partition of $G$ weighted by the spanning tree weight is balanced with probability at least ${1}/{q_{k}(|V|)}$.
Proof
@@
We note that the probability that a random spanning tree is $k$-splittable is the ratio of the number of $k$-splittable spanning trees to the number of spanning trees of a graph, and that the probability that a random $k$-partition weighted by the spanning tree weight is balanced is the ratio of the total weight of all balanced $k$-partitions to the total weight of all $k$-partitions. Then as a direct consequence of the next two lemmas, the ratio of the probability that a random spanning tree of $G$ is $k$-splittable to the probability that a random spanning tree weighted $k$-partition of $G$ is balanced is between $n^{-2(k+1)}$ and $n^{3(k+1)}$.
Lemma 3
@@
For any connected graph $G$ with $n$ vertices, the ratio between the total weight of all $k$-partitions of $G$ and the number of spanning trees of $G$ is between $n^{-2(k-1)}$ and $n^{k-1}$.
Proof
@@
The weight of a partition $P$ under the spanning tree weight is the product of the number of spanning trees in each piece. Then this weight is equal to the number of spanning forests whose connected components are the pieces of $P$.
@@
Consider the space of spanning forests in $G$ with $k$ connected components. We see that each such forest can, by adding k1k-1 edges of the components, be made into a spanning tree, and since every spanning tree can be split into $k$ pieces by arbitrarily removing k1k-1 edges, every spanning tree can be made this way. Since there are at most $n^{2}$ edges in $G$, the number of spanning trees is at most $n^{2(k-1)}$ times the number of such forests. Similarly, since every such forest can be made by removing k1k-1 edges from a spanning tree, and each tree contains $\binom{n-1}{k-1}$ forests with $k$ connected components, the number of such forests is at most $\binom{n-1}{k-1}<n^{k-1}$ times the number of spanning trees.
Lemma 4
@@
For any connected graph $G=(V,E)$ with $n$ vertices, the ratio of the total weight of all balanced $k$-partitions of $G$ and the number of $k$-splittable spanning trees of $G$ is between $n^{-2(k-1)}$ and $1$.
Proof
@@
By Lemma 1, each $k$-splittable spanning tree uniquely defines a $k$-balanced spanning forest, so the number of them is at most the number of spanning trees. Further, each $k$-splittable tree can be defined by a balanced forest and k1k-1 edges, so there are at most $\binom{|E|}{k-1}<n^{2(k-1)}$ times as many spanning trees as balanced forests.
2.2 Algorithms for generating balanced forests
@@
It will be useful to formalize the three main algorithms for sampling balanced partitions. First, the simplest algorithm is splitting a full spanning tree (see Algorithm 1). This algorithm will run a single stage and either return a balanced partition sampled uniformly according to the spanning tree weight, or fail. When using it in practice, the algorithm will be run until it succeeds in finding a balanced $k$-partition. It is described in [9]. The algorithm first generates a random spanning tree. Then it generates a partition by splitting the tree into $k$ equal pieces, if that is possible. Then it accepts the partition with probability equal to the inverse of the number of spanning trees that would generate that partition. The algorithm finds k1k-1 edges in the spanning tree to split it equitably as follows. If k1k-1 such edges exist, they are unique, according to Lemma 1. Each edge, when removed by itself from the tree, splits the tree into two pieces whose sizes are multiples of $n$, and this property uniquely characterizes these k1k-1 edges.
@@
6:  remove the edges in $K$ from the tree to create a forest;
@@
Next, we will examine two Markov-chain based algorithms. The up-down forest walk, described in [12], moves among forests with NkN-k edges (see Algorithm 2). The number of times $t$ that the loop is iterated depends on the mixing time, which was shown to be $O(n^{2}\log(n))$ in the worst case [12].
@@
The final algorithm we describe is Reversible ReCom as described in [8] (see Algorithm 3). This is also a Markov Chain algorithm that only moves between balanced partitions, so always returns a valid balanced partition, however its mixing time is unknown and sometimes exponential.
@@
All of the listed Markov Chain algorithms are known to converge to a stationary distribution that selects each forest with equal probability, including both reversible ReCom and Forest ReCom[6, 8, 12]. However, ReCom Markov chains may not be irreducible [3], so may have multiple stationary distributions depending on the starting partition. Moreover, guarantees on the mixing times of ReCom are unknown, with some cases known to require exponential time[12].
3 Obstacles, even with slack